Science Report Nov 15th

Science Report 15 November 2019

SOL: 5

Crew 215 – Expedition Boomerang Crew

Submitted by Science Officer Andrew Wheeler

Science Dome Operations: More micrometeorite samples were collected in the morning and

processed in the afternoon. Further gypsum samples were ground in preparation for the

cement making experiments to be carried out later in the rotation.

Green Hab: Green Hab operations have continued. See Green Hab report.

RAM Operations: N/A

EVA: The fifth EVA for science saw us returning to the MM5 (518462E 4248949N) site

where 24 micrometeorite samples were collected. 64% of the grid has now been sampled.

See EVA Report.

EVA Suit Maintenance: Nominal. See Operations Report.

Psycho-social Stressors Monitoring: Data collection continues through the hexoskin

monitoring vests. The simulated emergency retrieval of the engineer was carried out

during the engineering EVA and the performance of all participants is being evaluated.

In addition, indicated modifications to the airlock that would make this activity

easier are being evaluated.

Additional activities: N/A.

Copyright © The Mars Society. All rights reserved. | Main Site